# A General Latent Embedding Approach for Modeling Non-uniform High-dimensional Sparse Hypergraphs with Multiplicity

### v1

@@ 1 Introduction
We denote a hypergraph by ${\cal H}({\cal V}_{n},{\cal E}_{m})$, where ${\cal V}_{n}=\{v_{1},v_{2},\cdots,v_{n}\}$ represents the collection of $n$ vertices (individuals) and ${\cal E}_{m}=\{e_{1},e_{2},\cdots,e_{m}\}$ represents the collection of $m$ hyperlinks/hyper-edges (interactions). For notation simplicity, we denote $[n]=\{1,\cdots,n\}$ and ${\cal V}_{n}=[n]$. Each hyperlink $e_{j}$, for $j\in[m]$, consists of a subset of vertices that form the $j$-th hyperlink; that is, $e_{j}\subseteq[n]$ with $i\in e_{j}$ if and only if vertex $i$ is part of the $j$-th hyperlink. For example, if vertices $1,2$ and $3$ form the first hyperlink, we have $e_{1}=\{1,2,3\}$. A hyperlink consisting of $d$ vertices is termed an order-$d$ hyperlink. If all hyperlinks $e_{j}\in{\cal E}_{m}$ have the same order $d$, the hypergraph ${\cal H}({\cal V}_{n},{\cal E}_{m})$ is termed a $d$-uniform hypergraph. For instance, a typical graph network is essentially a 2-uniform hypergraph without repetitive hyperlinks. Conversely, if the hyperlinks in ${\cal E}_{m}$ vary in their orders, the hypergraph is described as non-uniform. Non-uniform hypergraphs are ubiquitous in the real world. For example, consider co-authorship relationships, where authors are vertices and each paper is treated as a hyperlink that includes its co-authors. The number of co-authors for each paper, i.e., the order of the hyperlink, naturally varies across different papers. Furthermore, it is common for the same set of co-authors to collaborate on multiple papers, leading to identical hyperlinks appearing multiple times within a hypergraph. This phenomenon is referred to as the multiplicity of hyperlinks. The non-uniformity and multiplicity of hyperlinks pose unique challenges in modeling hypergraphs.
@@
where $p_{ji}=g\big{(}\beta_{m,n}+\alpha_{i}+{\cal K}(\mathbf{f}_{j},\mathbf{z}_{i})\big{)}\in$ $(0,1)$ and $\boldsymbol{\Xi}_{m,n}=\big{\{}\beta_{m,n},\{\alpha_{i}\}_{i=1}^{n},\{\mathbf{z}_{i}\}_{i=1}^{n},\{\mathbf{f}_{j}\}_{j=1}^{m}\big{\}}$. Here, ${\cal K}(\cdot,\cdot):\mathbb{R}^{K}\times\mathbb{R}^{K}\to\mathbb{R}$ measures the “similarity” (to be defined) between the two latent embeddings, and $g(\cdot):\mathbb{R}\to[0,1]$, a monotonically increasing function, maps a real value to a probability. The $\{\alpha_{i}\}_{i\in[n]}$ parameters (with $\alpha_{i}\in\mathbb{R}$) account for heterogeneity in vertex degrees, which can be interpreted as a popularity measure of the vertices, and $\beta_{m,n}\in\mathbb{R}$, the order-adjusting parameter, controls the overall order of hyperlinks. Unlike most existing hypergraph modeling approaches where the order of hyperlinks is uniform, our model accommodates varying hyperlink orders. However, the orders of hyperlinks typically remain much smaller than both the number of vertices within the hypergraph and the number of observed hyperlinks, which poses an additional modeling challenge. $\beta_{m,n}$ serves to address this challenge by monitoring the overall order of hyperlinks compared to $n$ and $m$. In model (1.1), $\{\mathbf{z}_{i}\}_{i\in[n]}$ and $\{\mathbf{f}_{j}\}_{j\in[m]}$ can be treated either as realizations of some embedding distribution or as parameters to estimate. We treat them as fixed parameters in this work. In situations where these embeddings are treated as random vectors, our analysis can be regarded as making statistical inferences conditional on them. Moreover, it is worth noting that different $E_{j}$’s have different distributions. This means that rather than assuming an identical generating mechanism for all $m$ hyperlinks, our framework permits individualized probability assignments for each observed hyperlink, which is more flexible in modeling real-world hypergraphs. More discussion and interpretation of (1.1) are in Section 2.
@@ 1.1 Related work
Recent research has seen a surge in the direct modeling and analysis of hypergraphs. For instance, Ke et al. (2019) proposed a regularized tensor power iteration algorithm called Tensor-SCORE for community detection on uniform hypergraphs. Extensions to non-uniform hypergraphs were suggested without further investigation. The hypergraph degree-corrected block model introduced by Ke et al. (2019) incorporates vertex degree heterogeneity into the framework of the hypergraph planted partition model (Ghoshdastidar & Dukkipati, 2017). Yuan & Qu (2021) developed a tensor-based joint network embedding method that encodes both pairwise links and a subgroup of uniform hyperlinks into a latent space, capturing the dependency between pairwise and multi-way links in predicting potential unobserved hyperlinks. Yuan et al. (2022) conducted an extensive study on hypergraph testing for uniform hypergraphs under a spectrum of hyperlink probability scenarios, providing comprehensive theoretical analysis and significant insights into the subject. Zhen & Wang (2022) introduced a null vertex to augment a non-uniform hypergraph into a uniform hypergraph for community detection purposes. However, most existing hypergraph modeling approaches have significant limitations, including a heavy reliance on uniform hypergraph structures and the inability to account for repeated observations of identical hyperlinks due to the use of adjacency tensor representation. In the co-authorship data example, this means that only combinations of authors of a fixed size can be effectively modeled, and information about the number of papers a group of authors has co-authored is lost. In reality, the number of co-authors can vary for different papers, and the same set of co-authors can write more than one paper. Furthermore, the handling of non-uniform hypergraphs remains ad hoc. There is a pressing demand for a general modeling framework.
@@ 1.2 Contributions
Second, we investigate the identifiability conditions for the embeddings and associated parameters, and establish convergence rates for their estimators along with their asymptotic distributions. In Section 3.1, we introduce estimators that achieve $F$-consistency for the “${\boldsymbol{\Theta}}$” parameter (to be defined), as well as for the vertex popularity parameters and the order-adjusting parameter. In Section 3.2, we address the challenge of estimating individual components in $\boldsymbol{\Xi}_{m,n}$. Additional identifiability conditions are necessary to disentangle individual parameters, and we design a penalty term based on these conditions. Notably, this penalty term is not for shrinkage or regularization purposes, but for the implementation of a Lagrangian method to address identifiability issues. Furthermore, we allow the order-adjusting parameter $\beta_{m,n}$ to diverge to $\infty$- ∞ to accommodate the sparsity of the order of hyperlinks. This divergence results in irregular behavior of the likelihood function. We dedicate substantial effort to understanding this effect. Given that $\beta_{m,n}$ is designed to control the rate of hyperlink orders in the model, which will be further illustrated in Section 2, this exploration offers valuable insights into the types of datasets that are well-suited to our methodology. As a result, we establish entry-wise consistent estimators for individual parameters in $\boldsymbol{\Xi}_{m,n}$ and characterize their asymptotic distributions. These theoretical results subsequently permit estimation and inference on the probability terms $p_{ji}$’s.
@@ 1.3 Organization of the paper and notations
Throughout this work, we use regular letters, bold regular letters, and bold capital letters to denote scalars, vectors, and matrices, respectively. For any $a,b\in\mathbb{R}$, we denote their maximum and minimum by $a\vee b=\max(a,b)$ and $a\wedge b=\min(a,b)$, respectively. Given any two sequences $(a_{m,n}),(b_{m,n})$ valued in $\mathbb{R}$, we say $a_{m,n}\lesssim b_{m,n}$ (or $a_{m,n}=O(b_{m,n})$, $b_{m,n}\gtrsim a_{m,n}$) if there exists a universal constant $C>0$ such that $|a_{m,n}|\leq C|b_{m,n}|$. We say $a_{m,n}\asymp b_{m,n}$ if $a_{m,n}\lesssim b_{m,n}$ and $a_{m,n}\gtrsim b_{m,n}$. We write $a_{m,n}\ll b_{m,n}$ (or $a_{m,n}=o(b_{m,n})$) if $\lim_{m,n\to\infty}a_{m,n}/b_{m,n}=0$ and $a_{m,n}\gg b_{m,n}$ (or $a_{m,n}=\Omega(b_{m,n})$) if $\lim_{m,n\to\infty}a_{m,n}/b_{m,n}=\infty$. We use $O_{p}(\cdot)$, $\Omega_{p}(\cdot)$, and $o_{p}(\cdot)$ as the stochastic versions of $O(\cdot)$, $\Omega(\cdot)$, and $o(\cdot)$ for convergence in probability. For any positive integer $a$, we use $[a]$ to denote the set $\{1,2,\ldots,a\}$. For any column vector $\mathbf{a}$ and matrix $\mathbf{A}$, we use $\mathbf{a}^{\top}$ and $\mathbf{A}^{\top}$ to denote their transposes, respectively. For an event $\mathfrak{E}$, the indicator function $\mathbf{1}_{\mathfrak{E}}=1$ if $\mathfrak{E}$ occurs and 00 otherwise. For a vector $\mathbf{a}=(a_{1},\cdots,a_{n})^{\top}$, we use $\|\mathbf{a}\|_{2}$ to denote the standard Euclidean norm, $\|\mathbf{a}\|_{0}=\sum_{i\in[n]}\mathbf{1}_{\{a_{i}\neq 0\}}$ to denote the $L_{0}$ norm (the number of non-zeros), and $\|\mathbf{a}\|_{\infty}=\max_{i}|a_{i}|$ to denote the infinity/maximum/supreme norm. For a matrix $\mathbf{A}=(a_{ji})_{m\times n}$, we use $\|\mathbf{A}\|_{F}$ and $\|\mathbf{A}\|_{2}$ to denote its Frobenius and spectral norms, respectively. We use $\|\mathbf{A}\|_{\max}=\max_{j\in[m],i\in[n]}|a_{ji}|$ to denote the matrix maximum/supreme norm. For a square matrix $\mathbf{A}$, we use $\rho_{\max}(\mathbf{A})$ and $\rho_{\min}(\mathbf{A})$ to denote its maximal and minimal eigenvalues, respectively.
@@ 2 The order adjusted latent embedding approach
for some constant $C>0$. We allow $\beta_{m,n}$ to diverge to $\infty$- ∞ as $(m,n)\to\infty$, thus $p_{ji}$’s can be arbitrarily close to 00 in the asymptotic regime. Consequently, the expected order of hyperlinks (i.e. the expected number of vertices in a hyperlink) can be sublinear with respect to the size of the hypergraph. This is consistent with the observation in practice that the order of hyperlinks often increases at a much slower rate compared to the growth of the number of vertices.
To further understand (2.1), consider a specific hyperlink indexed by $j\in[m]$. For a subset of vertices $e\subset[n]$ with $|{e}|=d$, given $\boldsymbol{\Xi}_{m,n}$, model (1.1) along with (2.1) specifies the conditional probability that these vertices constitute hyperlink $j$ as
@@
The denominator of this probability remains constant regardless of the specific subset of vertices ${e}$. The numerator depends on a combination of factors: the order-adjusting parameter $\beta_{m,n}$, the vertices’ average popularity $\sum_{i\in{e}}\alpha_{i}/d$, the inner product between the hyperlink’s embedding $\mathbf{f}_{j}$ and the centroid of the vertices’ embeddings $\sum_{i\in{e}}\mathbf{z}_{i}/d$, and the order of the hyperlink $d$. The multiplicative operations in the second line of (2.3) suggest a form of conditional independence among the vertices given the embeddings. Such conditional independence has been widely considered in (hyper)graph network models, for example, by Holland et al. (1983), Hoff et al. (2002), Ke et al. (2019), and Ma et al. (2020). In this work, we do not model the distributions of embeddings as they can be general. Below,
@@
  • Average popularity of vertices: In Figure 1 (b), $\mathbf{f}_{j}^{\top}\boldsymbol{\mu}_{\mathbf{z},1}$ and $\mathbf{f}_{j}^{\top}\boldsymbol{\mu}_{\mathbf{z},3}$ are close, and the two groups have the same number of vertices. We consider the term $\sum_{i\in{e}}\alpha_{i}/d$, which we refer to as the average popularity of vertices in ${e}$. Note that larger dot sizes correspond to higher $\alpha$ values in Figure 1. Due to their higher average popularity, the group of vertices centered around $\boldsymbol{\mu}_{\mathbf{z},1}$ has a higher probability of forming hyperlink $j$ than the group centered around $\boldsymbol{\mu}_{\mathbf{z},3}$.

### v2

@@ 1 Introduction
We denote a hypergraph by ${\cal H}({\cal V}_{n},{\cal E}_{m})$, where ${\cal V}_{n}=\{v_{1},v_{2},\cdots,v_{n}\}$ represents the collection of $n$ vertices (individuals) and ${\cal E}_{m}=\{e_{1},e_{2},\cdots,e_{m}\}$ represents the collection of $m$ hyperlinks/hyper-edges (interactions). For notation simplicity, we denote $n]=\{1,\cdots,n\}$ and ${\cal V}_{n}=[n]$. Each hyperlink $e_{j}$, for $j\in[m]$, consists of a subset of vertices that form the $j$-th hyperlink; that is, $e_{j}\subseteq[n]$ with $i\in e_{j}$ if and only if vertex $i$ is part of the $j$-th hyperlink. For example, if vertices $1,2$ and $3$ form the first hyperlink, we have $e_{1}=\{1,2,3\}$. A hyperlink consisting of $d$ vertices is termed an order-$d$ hyperlink. If all hyperlinks $e_{j}\in{\cal E}_{m}$ have the same order $d$, the hypergraph ${\cal H}({\cal V}_{n},{\cal E}_{m})$ is termed a $d$-uniform hypergraph. For instance, a typical graph network is essentially a 2-uniform hypergraph without repetitive hyperlinks. Conversely, if the hyperlinks in ${\cal E}_{m}$ vary in their orders, the hypergraph is described as non-uniform. Non-uniform hypergraphs are ubiquitous in the real world. For example, consider co-authorship relationships, where authors are vertices and each paper is treated as a hyperlink that includes its co-authors. The number of co-authors for each paper, i.e., the order of the hyperlink, naturally varies across different papers. Furthermore, it is common for the same set of co-authors to collaborate on multiple papers, leading to identical hyperlinks appearing multiple times within a hypergraph. This phenomenon is referred to as the multiplicity of hyperlinks. The non-uniformity and multiplicity of hyperlinks pose unique challenges in modeling hypergraphs.
@@
where $p_{ji}=g\big(\beta_{m,n}+\alpha_{i}+{\cal K}(\mathbf{f}_{j},\mathbf{z}_{i})\big)\in$ $(0,1)$ and $\boldsymbol{\Xi}_{m,n}=\big\{\beta_{m,n},\{\alpha_{i}\}_{i=1}^{n},\{\mathbf{z}_{i}\}_{i=1}^{n},\{\mathbf{f}_{j}\}_{j=1}^{m}\big\}$. Here, ${\cal K}(\cdot,\cdot):\mathbb{R}^{K}\times\mathbb{R}^{K}\to\mathbb{R}$ measures the “interaction” (to be defined) between the two latent embeddings, and $g(\cdot):\mathbb{R}\to[0,1]$, a monotonically increasing function, maps a real value to a probability. The $\{\alpha_{i}\}_{i\in[n]}$ parameters (with $\alpha_{i}\in\mathbb{R}$) account for heterogeneity in vertex degrees, which can be interpreted as a popularity measure of the vertices, and $\beta_{m,n}\in\mathbb{R}$, the order-adjusting parameter, controls the overall order of hyperlinks. Unlike most existing hypergraph modeling approaches where the order of hyperlinks is uniform, our model accommodates varying hyperlink orders. However, the orders of hyperlinks typically remain much smaller than both the number of vertices within the hypergraph and the number of observed hyperlinks, which poses an additional modeling challenge. $\beta_{m,n}$ serves to address this challenge by monitoring the overall order of hyperlinks compared to $n$ and $m$. In model (1.1), $\{\mathbf{z}_{i}\}_{i\in[n]}$ and $\{\mathbf{f}_{j}\}_{j\in[m]}$ can be treated either as realizations of some embedding distribution or as parameters to estimate. We treat them as fixed parameters in this work. In situations where these embeddings are treated as random vectors, our analysis can be regarded as making statistical inferences conditional on them. Moreover, it is worth noting that different $E_{j}$’s have different distributions. This means that rather than assuming an identical generating mechanism for all $m$ hyperlinks, our framework permits individualized probability assignments for each observed hyperlink, which is more flexible in modeling real-world hypergraphs. More discussion and interpretation of (1.1) are in Section 2.
@@ 1.1 Related work
Recent research has seen a surge in the direct modeling and analysis of hypergraphs. For instance, Stasi et al. (2014) introduced the hypergraph $\beta$-model to capture node-degree heterogeneity in hypergraph networks. Nandy & Bhattacharya (2024) subsequently conducted a rigorous theoretical study of the hypergraph $\beta$-model, establishing convergence rates for its maximum likelihood estimator and detection thresholds for likelihood ratio tests with their optimality demonstrated (Chatterjee et al., 2011; Mukherjee et al., 2018). Ke et al. (2019) proposed a regularized tensor power iteration algorithm, called Tensor-SCORE, for community detection on uniform hypergraphs. Extensions to non-uniform hypergraphs were suggested without further investigation. The hypergraph degree-corrected block model introduced by Ke et al. (2019) incorporates vertex degree heterogeneity into the framework of the hypergraph planted partition model (Ghoshdastidar & Dukkipati, 2017). Yuan & Qu (2021) developed a tensor-based joint network embedding method that encodes both pairwise links and a subgroup of uniform hyperlinks into a latent space, capturing the dependency between pairwise and multi-way links in predicting potential unobserved hyperlinks. Yuan et al. (2022) conducted an extensive study on hypergraph testing for uniform hypergraphs under a spectrum of hyperlink probability scenarios, providing comprehensive theoretical analysis and significant insights into the subject. Zhen & Wang (2022) introduced a null vertex to augment a non-uniform hypergraph into a uniform hypergraph for community detection purposes. However, most existing hypergraph modeling approaches have significant limitations, including a heavy reliance on uniform hypergraph structures and the inability to account for repeated observations of identical hyperlinks due to the use of adjacency tensor representation. In the co-authorship data example, this means that only combinations of authors of a fixed size can be effectively modeled, and information about the number of papers a group of authors has co-authored is lost. In reality, the number of co-authors can vary for different papers, and the same set of co-authors can write more than one paper. Furthermore, the handling of non-uniform hypergraphs remains ad hoc. A recent work Turnbull et al. (2024) introduced a Bayesian latent space hypergraph model with a posterior sampling scheme for non-uniform hypergraphs, which does not scale to the high-dimensional hypergraphs considered in our setting and cannot handle hyperlink multiplicity. In our view, there is a pressing demand for a general modeling framework for high-dimensional non-uniform hypergraphs with multiplicity.
@@ 1.2 Contributions
Second, we investigate the identifiability conditions for the embeddings and associated parameters, and establish convergence rates for their estimators along with their asymptotic distributions. In Section 3.1, we introduce estimators that achieve $F$-consistency for the $\boldsymbol{\Theta}$}” parameter (to be defined), as well as for the vertex popularity parameters and the order-adjusting parameter. In Section 3.2, we address the challenge of estimating individual components in $\boldsymbol{\Xi}_{m,n}$. Additional identifiability conditions are necessary to disentangle individual parameters, and we design a penalty term based on these conditions. Notably, this penalty term is not for shrinkage or regularization purposes, but for the implementation of a Lagrangian method to address identifiability issues. Furthermore, we allow the order-adjusting parameter $\beta_{m,n}$ to diverge to -$\infty$ to accommodate the sparsity of the order of hyperlinks. This divergence results in irregular behavior of the likelihood function. We dedicate substantial effort to understanding this effect. Given that $\beta_{m,n}$ is designed to control the rate of hyperlink orders in the model, which will be further illustrated in Section 2, this exploration offers valuable insights into the types of datasets that are well-suited to our methodology. As a result, we establish entry-wise consistent estimators for individual parameters in $\boldsymbol{\Xi}_{m,n}$ and characterize their asymptotic distributions. These theoretical results subsequently permit estimation and inference on the probability terms $p_{ji}$’s.
@@ 1.3 Organization of the paper and notations
Throughout this work, we use regular letters, bold regular letters, and bold capital letters to denote scalars, vectors, and matrices, respectively. For any $a,b\in\mathbb{R}$, we denote their maximum and minimum by $a\vee b=\max(a,b)$ and $a\wedge b=\min(a,b)$, respectively. Given any two sequences $(a_{m,n}),(b_{m,n})$ valued in $\mathbb{R}$, we say $a_{m,n}\lesssim b_{m,n}$ (or $a_{m,n}=O(b_{m,n})$, $b_{m,n}\gtrsim a_{m,n}$) if there exists a universal constant $C>0$ such that $|a_{m,n}|\leq C|b_{m,n}|$. We say $a_{m,n}\asymp b_{m,n}$ if $a_{m,n}\lesssim b_{m,n}$ and $a_{m,n}\gtrsim b_{m,n}$. We write $a_{m,n}\ll b_{m,n}$ (or $a_{m,n}=o(b_{m,n})$) if $\lim_{m,n\to\infty}a_{m,n}/b_{m,n}=0$ and $a_{m,n}\gg b_{m,n}$ (or $a_{m,n}=\Omega(b_{m,n})$) if $\lim_{m,n\to\infty}a_{m,n}/b_{m,n}=\infty$. We use $O_{p}(\cdot)$, $\Omega_{p}(\cdot)$, and $o_{p}(\cdot)$ as the stochastic versions of $O(\cdot)$, $\Omega(\cdot)$, and $o(\cdot)$ for convergence in probability. For any positive integer $a$, we use $[a]$ to denote the set $\{1,2,\ldots,a\}$. For any column vector $\mathbf{a}$ and matrix $\mathbf{A}$, we use $\mathbf{a}^{\top}$ and $\mathbf{A}^{\top}$ to denote their transposes, respectively. For an event $\mathfrak{E}$, the indicator function $\mathbf{1}_{\mathfrak{E}}=1$ if $\mathfrak{E}$ occurs and 0 otherwise. For a vector $\mathbf{a}=(a_{1},\cdots,a_{n})^{\top}$, we use $\|\mathbf{a}\|_{2}$ to denote the standard Euclidean norm, $\|\mathbf{a}\|_{0}=\sum_{i\in[n]}\mathbf{1}_{\{a_{i}\neq 0\}}$ to denote the $L_{0}$ norm (the number of non-zeros), and $\|\mathbf{a}\|_{\infty}=\max_{i}|a_{i}|$ to denote the infinity/maximum/supreme norm. For a matrix $\mathbf{A}=(a_{ji})_{m\times n}$, we use $\|\mathbf{A}\|_{F}$ and $\|\mathbf{A}\|_{2}$ to denote its Frobenius and spectral norms, respectively. We use $\|\mathbf{A}\|_{\max}=\max_{j\in[m],i\in[n]}|a_{ji}|$ to denote the matrix maximum/supreme norm. For a square matrix $\mathbf{A}$, we use $\rho_{\max}(\mathbf{A})$ and $\rho_{\min}(\mathbf{A})$ to denote its maximal and minimal eigenvalues, respectively.
@@ 2 The order adjusted latent embedding approach
for some constant $C>0$. We allow $\beta_{m,n}$ to diverge to -$\infty$ as $(m,n)\to\infty$, thus $p_{ji}$’s can be arbitrarily close to 0 in the asymptotic regime. Consequently, the expected order of hyperlinks (i.e. the expected number of vertices in a hyperlink) can be sublinear with respect to the size of the hypergraph. This is consistent with the observation in practice that the order of hyperlinks often increases at a much slower rate compared to the growth of the number of vertices.
To further understand (2.1), consider a specific hyperlink indexed by $j\in[m]$. For a subset of vertices $e\subset[n]$ with |e|=d|{e}|=d, given $\boldsymbol{\Xi}_{m,n}$, model (1.1) along with (2.1) specifies the conditional probability that these vertices constitute hyperlink $j$ as
@@
The denominator of this probability remains constant regardless of the specific subset of vertices e{e}. The numerator depends on a combination of factors: the order-adjusting parameter $\beta_{m,n}$, the vertices’ average popularity $\sum_{i\in{e}}\alpha_{i}/d$, the inner product between the hyperlink’s embedding $\mathbf{f}_{j}$ and the centroid of the vertices’ embeddings $\sum_{i\in{e}}\mathbf{z}_{i}/d$, and the order of the hyperlink $d$. The multiplicative operations in the second line of (2.3) suggest a form of conditional independence among the vertices given the embeddings. Such conditional independence has been widely considered in (hyper)graph network models, for example, by Holland et al. (1983), Hoff et al. (2002), Ke et al. (2019), and Ma et al. (2020). In this work, we do not model the distributions of embeddings as they can be general. Below,
@@
  • Average popularity of vertices: In Figure 1 (b), $\mathbf{f}_{j}^{\top}\boldsymbol{\mu}_{\mathbf{z},1}$ and $\mathbf{f}_{j}^{\top}\boldsymbol{\mu}_{\mathbf{z},3}$ are close, and the two groups have the same number of vertices. We consider the term $\sum_{i\in{e}}\alpha_{i}/d$, which we refer to as the average popularity of vertices in e{e}. Note that larger dot sizes correspond to higher $\alpha$ values in Figure 1. Due to their higher average popularity, the group of vertices centered around $\boldsymbol{\mu}_{\mathbf{z},1}$ has a higher probability of forming hyperlink $j$ than the group centered around $\boldsymbol{\mu}_{\mathbf{z},3}$.
@@
Remark 2.1.
Our model uses hyperlink-vertex interactions to model hypergraph networks, whereas existing latent space models typically focus on vertex-vertex interactions. This hyperlink-vertex interaction modeling strategy enables us to move beyond the traditional framework and address the challenges posed by non-uniformity and multiplicity in general hypergraphs, while still preserving several key properties of existing models using vertex-vertex interactions. We refer the reader to Section 1 of the supplementary material for a detailed discussion.